# Radiating black holes in general relativity need not be singular

### v1

@@ V Conclusions
In this work we have provided an example of black hole formed via gravitational collapse that evaporates via Hawking radiation without producing any singularity nor Cauchy horizon.
A crucial point is that no exotic physics was added by hand. There is no need to modify the dynamics of gravitational interaction or to invoke exotic forms of matter. The combination of the repulsive effect due either to the electromagnetic interaction (for the simple case analyzed here) or the rotation of spacetime (for a more relevant astrophysical scenario) and the violation of energy conditions due to semiclassical gravity can be enough to avoid the breakdown of the validity of the theory. A classical mechanism of repulsion can produce an inner horizon and a bounce of the collapsing matter. This avoids the formation of the singularity inside the matter distribution. This does not contradict any singularity theorem as a singularity and/or a Cauchy horizon still form in the other asymptotic region. On the other hand, the role of semiclassical gravity (which is negligible within the collapsing matter) can prevent the formation of the Cauchy horizon, and the singularity beyond it, by violating the energy conditions and allowing the trapping region to disappear.
@@
Note added:
The day before this manuscript was uploaded, the paper Gralla:2025gzl , where a related idea is discussed, appeared on arXiv. There are some differences between the two works. The present one focuses on the singularity problem and tangentially touches issues related to the end of evaporation and the information loss problem. On the other hand, Gralla:2025gzl mostly focuses on the end point of the evaporation and the information loss problem, and at the end it explains the implications for the singularity problem. Furthermore, the cases in Fig. 3(b) and Fig. 3(c) discussed in the present work were not considered in Gralla:2025gzl . However, the overlap of results is substantial.

### v2

@@ V Conclusions
In this work we have shown that radiating black holes in general relativity need not be singular.
In fact, we have classified all possible spacetimes formed via gravitational collapse that evaporate via Hawking radiation based on their causual structure. We have seen that some of the possible scenarios are perfectly regular and contain neither a Cauchy horizon nor a curvature singularity.
A crucial point is that no exotic physics was added by hand. There is no need to modify the dynamics of gravitational interaction or to invoke exotic forms of matter. The combination of the repulsive effect due either to the electromagnetic interaction (for the simple case analyzed here) or to the rotation of spacetime (for a more relevant astrophysical scenario) and the violation of energy conditions due to semiclassical gravity can be enough to avoid the breakdown of the validity of the theory. A classical mechanism of repulsion can produce an inner horizon and a bounce of the collapsing matter. This avoids the formation of the singularity inside the matter distribution. This does not contradict any singularity theorem as a singularity and/or a Cauchy horizon still forms in the other asymptotic region. On the other hand, the role of semiclassical gravity (which is negligible within the collapsing matter) can prevent the formation of the Cauchy horizon, and the singularity beyond it, by violating the energy conditions and allowing the trapping region to disappear.
@@ VI Acknowledgments
The day before this manuscript was uploaded, the paper [38], where a related idea is discussed, appeared on arXiv. There are some differences between the two works. The present one focuses on the singularity problem and tangentially touches issues related to the end of evaporation and the information loss problem. On the other hand, [38] mostly focuses on the end point of the evaporation and the information loss problem, and at the end it explains the implications for the singularity problem. Furthermore, the cases in Fig. 3(a)3(c) and 3(d) discussed in the present work were not considered in [38]. However, the overlap of results is substantial.